OpenStudy (anonymous):

d/dt(v(t)=\sqrt{32t ^{2}}

OpenStudy (anonymous):

are you looking for the anti derivative?

OpenStudy (anonymous):

power rule of derivative

OpenStudy (anonymous):

are you trying to differentiate \[\sqrt{32t^2}\] or are you trying to find the function whose derivative is \[\sqrt{32t^2}\]?

OpenStudy (anonymous):

finding the function

OpenStudy (anonymous):

in other words \[\int\sqrt{32t^2}dt\]

OpenStudy (anonymous):

\[\sqrt{32}=4\sqrt{2}\] so this is \[4\sqrt{2}\int\sqrt{t^2}dt\]

OpenStudy (anonymous):

or even \[4\sqrt{2}\int |t|dt\]

OpenStudy (anonymous):

do you know how to solve it?

OpenStudy (anonymous):

so if t > 0 you get \[2\sqrt{2}t^2\]

OpenStudy (anonymous):

and if t < 0 you get \[-2\sqrt{2}t^2\]

OpenStudy (anonymous):

you can write this in one line if you put \[2\sqrt{2}t \sqrt{t^2}\]

OpenStudy (anonymous):

is that the power rule ?

OpenStudy (anonymous):

THANK YOU SO MUCH